# HandDiffuse: Generative Controllers for Two-Hand Interactions via Diffusion Models

### v1

@@ 4 Method
The first part of the Interaction Loss is the dual-hand joint contact potential loss. In each frame, we first calculate a $25\times 25$ distance matrix $\mathbf{D}$ between the 21 left-hand joints and 21 right-hand joints with 4 more joints sampled in each palm. Since the interaction between the hands involves fine-grained movements, we should focus on situations where the finger joints are very close to each other. Inspired by the concept of Contact Potential Field[56], we further represent the distances using the elastic potential energy of a spring model and it can be described as
$\mathbf{P}otential=1/2\mathbf{K}\ relu(\mathbf{\tau}-\|\mathbf{D}_{pred}\|_{2})^{2},$ (1)
@@
where $\mathbf{K}$ represents the coefficient of elasticity and $\mathbf{\tau}$ denotes the distance threshold. The closer the distance between the finger joints, the larger the potential energy obtained, and the distance will be filtered out when it is over the threshold. Moreover, by considering the direction of the distance, we can avoid ambiguity caused by penetration, and the final contact potential loss is described as
@@
where (×)($\times$)( × ) means cross product. More is shown in Figure 5.
Figure 7: Qualitative comparison for different methods. For Unconstrained Generation, MDM[52] tend to generate motion with hands shifting away from each other, while InterGen[31] would generate motions with a mean pose. In terms of Trajectory Control, MDM yields satisfactory results, whereas InterGen generates motion without incorporating interaction.
During our testing process, we observed that MANO has strict requirements for the proportional lengths of hand bones. Therefore, the second part of the Interaction Loss is the Shape Loss. We ensure the rationality of hand shape by calculating the differences in bone lengths between the left and right hands, as well as the differences between the predicted shape and the ground truth (GT). This is described as follows:
@@
where $BL$ denotes the Euclidean distance of the bones.
4.3 Downstream Application
For the task of unconstrained generation and motion in-betweening, the key challenge lies in constructing the global information of the dynamic interaction process between the two hands. Therefore, we choose to use the global rep. In motion in-betweening task, we provide the ground truth for the previous five frames $\mathbf{x_{0}^{1:5}}$ and the last five frames $\mathbf{x_{0}^{-5:-1}}$ for the diffused motion $\mathbf{x_{t}^{1:N}}$.
For the task of trajectory control, the goal is to generate the reasonable local motion given the global rotation and translation of both hands. In this case, using the local rep allows the network to focus on the local motion during the interaction process. Furthermore, we provide $\dot{R}^{1:N}$ and $\dot{T}^{1:N}$ for the diffuse $\mathbf{x_{t}^{1:N}}$.
5 Experiment
In this section, we conduct several experiments to evaluate our dataset and method for dual-hand motion generation tasks. We particularly evaluate (1) the comparison of our method against previous state-of-the-art approaches, (2) the ability of our method to augment existing datasets, and (3) the effectiveness of our hand motion representation and Interaction Loss. The above evaluations are conducted specifically on the three tasks mentioned in Section 4.3. The models have been trained with $T=1000$ noising steps and a cosine noise schedule. All of them have been trained on a single NVIDIA GeForce RTX 3080 GPU for about two days.
Methods Unconstrained In-betweening Trajectory Control
@@ 5 Experiment
Real 0.050 12.034
MDM[52] 0.526 10.366 0.616 10.258 0.205 11.265
InterGen[31] 0.251 4.928 0.269 5.385 0.402 9.102
Ours 0.176 11.786 0.282 11.533 0.182 11.523
Table 2: Quantitative comparisons of different methods. We compare the FID and Diversity for different methods. ’normal-$\downarrow$’ indicates that results are better with lower metrics. Bold means the best result.
5.1 Comparisons of different methods
To the best of our knowledge, our proposed method is the first to specifically target hand-only interaction motion generation. To provide a comprehensive evaluation, we compare our approach with two recent state-of-the-art human motion generation methods, namely MDM [52] and InterGen [31]. We modify these methods by adapting their motion representation to enable hand motion generation. As depicted in Figure 7, both MDM and InterGen exhibit limitations when generating hand motion. They tend to produce motion sequences with severe joint penetration or unnatural interaction since they were not originally designed for dual-hand motion generation. In contrast, our proposed method showcases superior performance as demonstrated by the lowest Fréchet Inception Distance (FID) values, as shown in Table 2. The way we adopted to calculate FID is introduced in the Appendix.
Figure 8: Cross data Evaluation. Motion In-betweening task result on InterHand2.6M[36]. We sample two frames in InterHand2.6M as the start frame and end frame and generate the in-betweening motion. Our HandDiffse can help augment the existing datasets.
5.2 Cross Datasets Evaluation
To evaluate our method cross datasets, we conduct experiments on two large-scale dual-hand interaction motion datasets, InterHand2.6M[36] and Re:InterHand [37] by sampling control signal from these two datasets. As shown in Figure 8, given the first frame and last frame from InterHand2.6M, our model is able to generate diverse motions. Furthermore, Table 3 shows that after performing In-betweening and Trajectory Control tasks on the datasets, the generated data can increase the diversity. This demonstrates the effectiveness of our approach and HandDiffuse can contribute to data augmentation for other datasets. More results and the details of experiments are introduced in the Appendix.
Datasets InterHand Re:InterHand
@@ 5.2 Cross Datasets Evaluation
Real 0.113 8.478 0.034 10.583
In-betweening 0.194 8.171 0.208 11.200
Root Control 0.142 10.146 0.236 12.979
Table 3: Cross Dataset Evaluation.The FID and Diversity both increase, indicating that we get a larger dataset from the original dataset by performing data generation tasks.
5.3 Ablation Study
As we propose two kinds of hand motion representations in section 4.1, here we demonstrate the effectiveness of the representations in different tasks by conducting qualitative and quantitative comparisons. Figure 9 illustrates the limitations of the model without Shape Loss and Interaction Loss. By incorporating the Interaction Loss, we mitigate the issue of joint penetration, resulting in improved motion generation quality. To provide a quantitative analysis, we employ FID and the penetration metric sdf loss(SDF)[23] for evaluation. SDF represents the Signed Distance Field, which could be defined as:
$\phi(x,y,z)=-\min(SDF(x,y,z),0),$ (4)
where $(x,y,z)$ denotes the position of MANO vertex. According to the definition, when the hand joints penetrate with each other, $\phi$ would take the positive value while zero otherwise. Using the formulation, we can measure the severity of penetration during the whole motion. Table 4 demonstrates that our model achieves the highest performance when using the proper motion representation and Interaction Loss.
Tasks Unconstrained In-betweening Trajectory Control
[…]
Real 0.050 1.146 - -
w Global rep. 0.168 1.575 0.282 0.415
w Local rep. 0.436 1.988 0.572 0.182
w/o Interaction Loss 0.176 1.652 0.318 0.216
Complete model 0.168 1.575 0.282 0.182
Table 4: Quantitative comparisons of different hand motion representations and w/o Interaction Loss. SDF indicates the sdf loss for penetration of joints during the whole motion. Note that the mean of sdf loss is only calculated on frames with penetraion.
6 Conclusion
Figure 9: Ablation Study. In the absence of Shape Loss, the model generates abnormal hand shapes. Similarly, the absence of Interaction Loss leads to hand interaction motion with severe penetration.
@@ 6 Conclusion
We present HandDiffuse12.5M, the largest interacting hands dataset with strong interaction and temporal sequences, to open up the research direction of interacting hands motion generation. HandDiffuse12.5M consists of various modalities, including synchronous 50-view videos, 3D keypoints and hand poses. Based on HandDiffuse12.5M, we propose a strong baseline method HandDiffuse which is the first interacting hands motion generation approach along with two specific controllers: key frames control and trajectory control. Our evaluation in multi-stage demonstrates the benefit of our approach againist others and HandDiffuse fills the blank in human motion generation.
Appendix
7 Experiment Details
7.1 Frechet Inception Distance (FID)
The Frechet Inception Distance (FID) is a metric used to quantify the dissimilarity between the distributions of generated and ground truth data in latent space. In our experiments, we employ FID as a crucial indicator to measure the authenticity and plausibility of our generated data. And the formula is
$FID=|\mu_{gt}-\mu_{gen}|^{2}+\text{Tr}(C_{gt}+C_{gen}-2\sqrt{C_{gt}C_{gen}}),$
where $\mu_{gt}$ and $\mu_{gen}$ are the mean feature vectors of ground truth and generated motions, respectively; $C_{gt}$ and $C_{gen}$ are the covariance feature matrices of real and generated motions, respectively; Tr()($\cdot$)( ⋅ ) denotes the trace of the matrix. In the previous works[51], the feature is extracted by a Inception network which is a classifier. However, in the interacting hands datasets, it is challenging to label these complex interactive movements, making it difficult to build a classifier. Therefore, taking inspiration from MotionCLIP[50], we propose a self-supervised encoder-decoder architecture for the reconstruction network, as depicted in Figure 10.
More specifically, we concatenate a $z_{tk}$ vector of equal dimensions to the original motion representation $x^{1:T}$. This concatenated sequence $x^{1:T+1}$ is then fed into a transformer encoder. From the encoder’s output, we select the corresponding position $z_{x}$ as the feature representation for the input motion segment. This obtained feature $z_{x}$ serves as the memory for the decoder. Additionally, we use a T-length zero vector that undergoes positional encoding as the target for reconstructing the motion in the decoder. The reconstruction loss is computed using the L2 norm.
During experimentation, we observed that using the global representation of the motion improves the convergence of the network. Therefore, in subsequent experiments, we convert the motion to a global representation before calculating the motion’s feature.
Figure 10: Feature Extractor for FID. Inspired by MotionCLIP[50], we train a transformer with encoder-decoder architecture by using self-supervised strategy.
7.2 Diversity
Diversity measures the variability in the resulting motion distribution. The way we followed to calculate Diversity is as same as MDM[51]. We randomly generate 1000 samples and select 200 of them for calculating.
8 HandDiffuse12.5M
Figure 11 has shown the diversity and complexity of our dataset HandDiffuse12.5M. Figure 12 has shown our manual annotation tool.
Figure 11: Overview of HandDiffuse12.5M. Each image represents one frame of hand motion from a different sequence.
Figure 12: Annotation Tool. Our specially designed tool for hand joint annotation of severely occluded frames.
9 More Results
Figure 13 has shown the results of trajectory control task. In each sequence, we adopt the wrists’ rotation and translation from InterHand2.6M[36] as our control signal and we can generate different motion based on this trajectory.
Figure 13: Results of Trajectory Control by Sampling from InterHand2.6M. Note that only Root orientation and position is given, the local motions for hand joints are generated from our model.

### v2

@@ Introduction
The first part of the $\mathbf{Loss}_{interaction}$ is the contact potential loss($\mathbf{Loss}_{potential}$ ). In each frame, we first calculate a $25\times 25$ distance matrix $\mathbf{D}$ between the 21 left-hand joints and 21 right-hand joints with 4 more joints sampled in each palm. Since the interaction between the hands involves fine-grained movements, we should focus on situations where the finger joints are very close to each other. Inspired by the concept of Contact Potential Field(Yang et al. 2021), we further represent the distances using the elastic potential energy of a spring model and it can be described as
$\mathbf{P}=1/2\mathbf{K}\ relu(\mathbf{\tau}-\|\mathbf{D}_{pred}\|_{2})^{2},$ (4)
@@
where $\mathbf{K}$ represents the coefficient of elasticity and $\mathbf{\tau}$ denotes the distance threshold. The closer the distance between the finger joints, the larger the potential energy obtained, and the distance will be filtered out when it is over the threshold.
Moreover, by considering the direction of the distance, we can avoid ambiguity caused by penetration, and the final contact potential loss is described as
@@
where (×)($\times$)( × ) means cross product. More is shown in Appendix.
Figure 7: Qualitative comparison for different methods. For Unconstrained Generation, MDM tend to generate motion with hands shifting away from each other, while InterGen would generate motions with a mean pose. In terms of Trajectory Control, MDM yields satisfactory results, whereas InterGen generates motion without incorporating interaction.
We observed that MANO has strict requirements for the bone length. Therefore, the second part of the $\mathbf{Loss}_{interaction}$ is the Shape Loss($\mathbf{Loss}_{shape}$ ). We ensure the rationality of hand shape by calculating the differences in bone lengths between the left and right hands, as well as the differences between the predicted shape and the ground truth (GT). This is described as follows:
@@
where $BL$ denotes the bone length.
Downstream Applications
For the task of key frames control, we provide the ground truth for the previous five frames $\mathbf{x_{GT}^{:5}}$ and the last five frames $\mathbf{x_{GT}^{-5:}}$ for the diffused motion $\mathbf{x_{t}^{1:N}}$. For the task of trajectory control, the goal is to generate the reasonable local motion given the global rotation and translation of both hands. In this case, we provide $\dot{R}^{1:N}$ and $\dot{T}^{1:N}$ as control signal for the diffused motion $\mathbf{x_{t}^{1:N}}$. The results are presented in Figure 6.
We also explore HandDiffuse by applying the control of 2D keypoints, and the experiments have shown that HandDiffuse can also solve the task of hands reconstruction. Due to the space limitation, the results are presented in Appendix.
Experiment
In this section, we conduct several experiments to evaluate our method HandDiffuse for the task of interacting hands motion generation. We particularly evaluate (1) the comparison of our method against previous state-of-the-art approaches, and (2) the ablation study. The above evaluations are conducted specifically on the two controllers (key frames and trajectory) and other controller(2D keypoints) is evaluated in Appendix due to the space limitation. More experiments on the ability of HandDiffuse to augment existing datasets are presented in Appendix. The models have been trained with $T=1000$ diffusing steps and a cosine noise schedule. The denoiser has 8 layers, 4 heads and the feedforward dimension is 1024. We set the length of generated motion $N$ to 200 in all experiments. All of them have been trained on a single NVIDIA GeForce RTX 3090 GPU for about two days.
Method Unconditioned In-betweening Trajectory Control
@@
real 0.050 12.034 1.146
MDM(Tevet et al. 2023) 0.526 10.366 1.74 0.616 10.258 1.621 0.205 11.265 1.572
InterGen(Liang et al. 2023) 0.251 4.928 0 0.269 5.385 1.032 0.402 9.102 3.572
BUDDI(Müller et al. 2024) 0.853 8.282 2.293 1.211 7.341 2.482 0.921 7.497 1.842
InterHandGen(Lee et al. 2024) 0.437 10.238 1.731 0.734 8.620 1.819 0.539 10.124 1.282
HandDiffuse(ours) 0.161 11.829 1.553 0.273 11.554 1.562 0.173 11.758 1.141
Table 2: Quantitative comparisons of different methods. We compare the FID and Diversity for different methods. $\downarrow$’ indicates that results are better with lower metrics, $\uparrow$ is on the contrary. Bold means the best result.
Task Unconditioned In-betweening Trajectory Control
@@
real 0.050 1.146 - - - -
w Global rep. 0.161 1.553 0.273 1.562 0.418 1.347
w Local rep. 0.431 1.962 0.567 1.971 0.173 1.141
w/o SHDe 0.168 1.575 0.282 1.571 0.182 1.538
w/o ${Loss}_{potential}$ 0.171 1.606 0.317 1.628 0.210 1.614
w/o ${Loss}_{shape}$ 0.390 1.583 0.512 1.579 0.380 1.245
Complete model 0.161 1.553 0.273 1.562 0.173 1.141
Table 3: Quantitative comparisons. “w/o SHDe” means IHDe is the only denoiser. SDF is only calculated on frames with penetraion.
Comparisons of different methods
We compare our approach with MDM (Tevet et al. 2023)(single human motion generation), InterGen (Liang et al. 2023)(interacting human motion generation), BUDDI(Müller et al. 2024)(latent representation of interacting human) and InterHandGen(Liang et al. 2023)(interacting hands motion generation of single frame). The metrics we applied are common used Fréchet Inception Distance (FID), Diversity and sdf loss(SDF). FID can measure the realistic of the generated motion. Diversity can prove that the motion do not maintain the mean pose for the sake of preserving realistic. SDF would take the positive value while zero otherwise, when the hand joints penetrate with each other. Using the formulation, we can measure the severity of penetration during the whole motion. The strategy we applied to compute these metrics is inspired by MotionCLIP(Tevet et al. 2022a) and is illustated in Appendix. We modify these methods by adapting their motion representation to enable hand motion generation. Specifically, we adapted InterHandGen to generate continuous frames following its original pipeline in which we first generated 200 frames of left-hand motion, then generated the right-hand motion based on the left hand. We adapted BUDDI for continuous frame generation and treat the motion of each frame as a token, and retained the ”Person Embed” component to distinguish different hands. As depicted in Figure 7, both MDM and InterGen exhibit limitations when generating hand motion. They tend to produce motion sequences with severe joint penetration, unnatural interaction or static mean pose. The resulst of HandDiffuse has the highest diversity as shown in Table 2. Because the motion generated by InterGen without condition is static, the FID is lower than us and the penetration does not exist.
Ablation Study
At first, we demonstrate the effectiveness of the different motion representations in different tasks in Table 3. The qualitative result is shown in Appendix. By incorporating the $\mathbf{Loss}_{interaction}$, we mitigate the issue of joint penetration, resulting in improved motion generation quality. Table 3 demonstrates that our Interacting Hands DDIM with modular deisgn achieves the highest performance when using the proper motion representation and $\mathbf{Loss}_{interaction}$.
Conclusion
@@
We present HandDiffuse12.5M, the largest interacting hands dataset with strong interaction and temporal sequences, to open up the research direction of interacting hands motion generation. We propose a strong baseline method HandDiffuse which is the first interacting hands motion generation approach along with some useful controllers: key frames control, trajectory control and 2D keypoints control. Our evaluation in multi-stage demonstrates the benefit of our approach againist others and HandDiffuse fills the blank in human motion generation.
Acknowledgments
We would like to express our sincere gratitude to all the students and faculty members of the VDI Laboratory at ShanghaiTech University for their invaluable assistance and guidance throughout this experiment. We also appreciate the use of the filming equipment provided by the VDI Laboratory.
Additionally, we would like to acknowledge our collaborators: Sihang Xu, Hongdi Yang, Yiran Liu, Xin Chen, Jingya Wang, Jingyi Yu, and Lan Xu, for their contributions to this work.